# Self-Supervised Learning in Event Sequences: A Comparative Study and Hybrid Approach of Generative Modeling and Contrastive Learning

### v1

@@ 5.2 Evaluation objectives
To evaluate the effectiveness of self-supervised training strategies, our study focuses on two key aspects: the quality of the embeddings and the effectiveness of fine-tuning the entire network, following self-supervised studies Dubois et al. (2024, 2021).
5.2.1 Main Objectives
To assess the quality of embeddings, we primarily rely on metrics from downstream tasks. This involves utilizing both linear and non-linear probing methods. For non-linear probing, we employ Gradient Boosting algorithms with LGBM Ke et al. (2017) applied directly to the embeddings.
We evaluate embeddings on several tasks, including the prediction of an attribute or a target $y_{i}$ given the entire $S_{i}$ and the prediction of consequent event $x^{j}$ given a set of $S_{i}=\{x_{i}^{1},\ldots,x^{j-1}_{i}\}$. The second objective addresses the TPP task, which involves predicting the type of the next event or the time of the next event’s arrival. We have processed each dataset such that the target is either the category of the next event (for datasets that include this feature) or the time until the next event (for datasets without a primary event category).
5.2.2 Secondary Objectives
We incorporate additional metrics to assess the quality of embeddings. Specifically, we utilize anisotropy and intrinsic dimension, drawing on insights from previous research Nakada and Imaizumi (2020); Razzhigaev et al. (2023). In Section 6.3, we compare our findings with the results presented in the aforementioned works.
Anisotropy assesses the non-uniform distribution of embeddings in space, providing insights into the contextualization of our embedding. Lower anisotropy in embeddings has been linked to better model performance Ait-Saada and Nadif (2023). In line with the approach used in Razzhigaev et al. (2023), we compute anisotropy as the ratio of the highest singular value to the sum of all singular values:
$Anisotropy(H)=\frac{\sigma_{1}^{2}}{\sum\limits_{i}\sigma_{i}^{2}}$
The intrinsic dimension evaluates the optimal dimensionality of data, shedding light on the core information captured by the embeddings. To determine the intrinsic dimension, we employ the method proposed in Facco et al. (2017). This method examines how the volume of an $n$-dimensional sphere changes with dimension $d$. Further details are available in the original paper or in Razzhigaev et al. (2023).
@@ 5.3 Models
Feature embeddings. To transform features into a numerical vector, we use an embedding layer for categorical features and linear projection for numerical ones. For ABank, Age and TaoBao we set the dimension of each feature to 32. For Pendulum and PhysioNet dimension is set to 8. Additionally, we use the time difference between events as a separate feature.
For consistency in our experiments, we use the same encoder architecture across all training strategies and datasets. We use a single-layer GRU with a hidden size of 512, we take the last hidden state as sequence embedding. The GRU is selected for its proven effectiveness in encoding time-ordered sequences and its extensive application in relevant literature, providing a reliable baseline for our study Rubanova et al. (2019); Tonekaboni et al. (2021); Yoon et al. (2019); Udovichenko et al. (2024).
@@
Contrastive learning approach is described at 4.1 and uses the same encoder as mentioned above. Furthermore, we integrate a Feed Forward Projector atop the GRU for loss calculation, a technique commonly adopted in several studies for enhanced performance Grill et al. (2020); Oquab et al. (2023).
@@ 5.4 Training details
To maintain consistency in our experiment, we trained all models using identical parameters. Models were trained for 100 epochs on datasets with fewer than 100K sequences and for 40 epochs on larger datasets. The learning rate was set at 1e-3, weight decay at 3e-3, and the batch size was fixed at 128. For the MLEM model, we set $\alpha$ at 1 and $\beta$ at 10. We averaged the results across three different seeds to ensure reliability and reproducibility.
@@ 6 Results
6.1 Self-supervised pre-training for fine-tuning
Here we evaluate the effectiveness of self-supervised pre-training strategies prior to supervised fine-tuning the entire model. For fine-tuning, we employed pre-trained encoders with randomly initialized linear layers and trained the model for 10 epochs. The results presented in Table 1 indicate that, among all the evaluated pre-training methods, MLEM consistently achieves the most favorable results across all datasets, with the exception of the highly sparse PhysioNet dataset. Furthermore, it is noteworthy that supervised learning without self-supervised pre-training consistently produces inferior results.
@@ 6.2 Embeddings probing
We conducted a comprehensive analysis of the quality of embeddings using both linear and non-linear probing techniques, results are presented in Table 2. We evaluated performance on various tasks described in Section 5.2.1, including EvS classification and regression tasks for the entire sequence, as well as TPP tasks. Despite some variations in absolute values, the overall trends and rankings of methods remain consistent across both probing strategies.
Our assessment reveals that neither the contrastive nor the generative approach consistently outperforms the other. However, the MLEM model consistently demonstrates superior performance across most datasets. Notably, when MLEM is not the top performer, it often ranks second, highlighting its versatility. By effectively integrating the advantages of both contrastive and generative techniques, this approach consistently delivers strong overall performance across a range of datasets.
Figure 4: Mean percentage change averaged across datasets. The X-axis represents the dropout probability, while the Y-axis indicates the mean change in the linear probing downstream metric, compared to the metric with no dropout.
6.3 Anisotropy and Intrinsic Dimension
@@
Table 3: Average anisotropy and intrinsic dimension across datasets
We conducted an evaluation of anisotropy and the intrinsic dimensions of embeddings from various pre-training strategies, as detailed in 5.2.2. As a result, we show that MLEM has the highest intrinsic dimension and the lowest anisotropy, on average, across all datasets. This finding may indicate that our pre-training strategy effectively enables the embeddings to encapsulate a greater amount of information and ensures their uniform distribution in space. The results are presented in Table 3.
Contrary to the findings in Dubois et al. (2024), we did not observe any correlation between intrinsic dimension and downstream performance metrics. Similarly, we did not find any correlation for anisotropy, which aligns with the aforementioned work.
Figure 5: Correlation plots for anisotropy and intrinsic dimension with normalized performance.
Figure 6: The t-SNE visualizations showcase embeddings resulting from various pre-training strategies. The top row displays the embeddings for the Age dataset, while the bottom row illustrates those for the TaoBao dataset. Each point represents a sequence $S_{i}$ from a given dataset, colored accordingly to the corresponding attribute $y_{i}$. For Age, there are 4 classes and 2 classes for TaoBao
6.4 Visualization
7 Embedding Robustness
Our study also explores the robustness of model embeddings to specific types of perturbations. We focus on examining the impact of dropout and the shuffling of events within one sequence on the performance of the embeddings on downstream tasks.
7.1 Sensitivity to data omission
To investigate the effects of data omission, we applied a random dropout strategy, removing entire events from a sequence along the time axis, with varying probabilities $p_{dropout}\in\{0.1,0.3,0.5,0.7\}$. This allowed us to examine the impact of different levels of data removal on the embeddings, as seen in Figure 4, by measuring the average percentage decrease in downstream task metrics.
MLEM’s performance slightly decreases at $0.1$ dropout probability and more so at higher levels. This indicates that MLEM is more sensitive to dropout and requires the presence of all the samples rather than depending on one specific sample. This property can be used in a sensitivity analysis or related applications.
Additionally, the generative modeling embeddings exhibited an intriguing behavior: they not only resisted the decline but actually surpassed their original performance at a dropout rate of 0.5. To draw reliable conclusions based on this finding, further work is required with a larger number of datasets and various models.
7.2 Sensitivity to perturbations
In subsequent tests, we examined the impact of perturbing events within a sequence on downstream task performance. The results, presented in Table 4, show that perturbing sequence elements had minimal to no effect on the results. This suggests that current models may be employing a ’Bag-of-Words’ approach and considering a set of events rather than the sequential nature of the data. One might assume that this discrepancy is due to not suitable models or training strategies. However, this hypothesis is not supported by a decline in performance on the pendulum dataset, where the time component is crucial and the same models were used.
This raises the question of whether we could further enhance model performance by explicitly considering the sequential nature of the data, or if this is solely a property of the datasets themselves.
@@ 7.2 Sensitivity to perturbations
Table 4: Percentage change in linear probing performance after shuffling events order
@@ 8 Computational complexity
To obtain the MLEM encoder, we need to train two models, which is not desirable. However, to demonstrate the practicality of our method, we compare MLEM with SeqNas Udovichenko et al. (2024) in terms of performance and computational efficiency. SeqNas employs an architectural search to identify the most optimal supervised architecture for the task at hand. We posit that SeqNas represents an upper bound for downstream metrics. Table 5 presents results from the two largest datasets, examined by us, demonstrating that MLEM achieves performance near the upper bound set by SeqNas, but with significantly lower computational cost.
@@
Table 5: Comparison of downstream metrics and computational costs for training and fine-tuning MLEM versus SeqNas.
9 Discussion
We observe that in our settings, MLEM outperforms the Naïve method, despite their initial similarities. One possible explanation for the success of our method can be attributed to viewing MLEM as a Knowledge Distillation Hinton et al. (2015) procedure. However, further work is required to confirm this hypothesis.
We observed that data perturbation does not significantly impact model performance, suggesting a near-time-invariance. This could be attributed to the nature of the examined datasets or the absence of appropriate modeling approaches. Both hypotheses suggest the potential for future research in time-sensitive modeling.
We also observe consistent underperformance of methods that include generative approach on the PhysioNet dataset. This could be attributed to the high rate of missing values, which complicates the generation of accurate predictions for subsequent steps. Furthermore, the relatively small size of the dataset may be a contributing factor, given that generative modeling is generally dependent on large amounts of data.
@@ 10 Conclusion
We proposed a novel approach to combine multiple self-supervised strategies and demonstrated that this approach yields superior results compared to any single self-supervised method. We believe this opens up a new direction for enhancing self-supervised learning.
Our study revealed that self-supervised pre-trained models, especially those with generative pre-training, outperform purely supervised methods, with MLEM showing the most promise. This suggests potential benefits in exploring different pre-training and architectural options.
In our work, we are the first, to our knowledge, to apply generative modeling to event sequences for representation learning and pre-training. Our findings show that neither generative nor contrastive pre-training clearly surpasses the other in effectiveness. To address this, we developed the MLEM method, which innovatively merges contrastive learning with generative modeling. MLEM aligns these two approaches as separate modalities, inspired by recent advances in multimodal research, and it demonstrates enhanced performance on most datasets for both primary and secondary goals.

### v3

@@ 4.2 Evaluation objectives
To evaluate the effectiveness of self-supervised training strategies, we follow previous studies Dubois et al. (2024, 2021) and focus on two key aspects:
  • Fine-Tuning Performance. The downstream metrics after fine-tuning the entire network.
  • Embeddings Probing. The quality of the embeddings for downstream and TPP tasks tested via linear and non-linear probing.
As well, we include additional objectives for embedding quality, that were discovered in previous studies Razzhigaev et al. (2023); Nakada and Imaizumi (2020) , that are connected with performance:
  • Anisotropy. It assesses the non-uniformity of embeddings in space, providing insights into the contextualization of our embedding. Lower anisotropy in embeddings has been linked to better model performance Ait-Saada and Nadif (2023).
  • Intrinsic dimension. It evaluates the optimal dimensionality of data, showing information captured by the embeddings.
Objectives details are described in appendix Objectives description.
@@ 4.3 Models
Base model and Feature embeddings. To transform features into a numerical vector, we use an embedding layer for categorical features and linear projection for numerical ones. For ABank, Age and TaoBao we set the dimension of each feature to 32. For Pendulum and PhysioNet dimension is set to 8. Additionally, we use the time difference between events as a separate feature.
We use GRU as an encoder because it has been proven effective in encoding time-ordered sequences  Rubanova et al. (2019); Tonekaboni et al. (2021); Yoon et al. (2019); Udovichenko et al. (2024). Udovichenko et al. performs a Neural Architecture Search on various models and highlights that the presence of RNN blocks in the architecture is associated with higher performance on average. Lin et al. illustrate that the performance gain brought by history encoders is very small for TPP modeling. Based on these empirical observations, we believe that our design choice is reasonable.
We it with a single-layer and a hidden size of 512, we take the last hidden state as sequence embedding. For consistency in our experiments, we use the same hyperparameters across all training strategies and datasets.
@@
Contrastive learning approach is described at 3.1 and uses the same encoder as mentioned above. Furthermore, we integrate a Feed Forward Projector atop the GRU for loss calculation, a technique commonly adopted in several studies for enhanced performance Grill et al. (2020); Oquab et al. (2023).
@@ 4.4 Training details
To maintain consistency in our experiment, we trained all models using identical parameters. Models were trained for 100 epochs on datasets with fewer than 100K sequences and for 40 epochs on larger datasets. The learning rate was set at 1e-3, weight decay at 3e-3, and the batch size was fixed at 128. For the Hybrid models, we set $\alpha$ at 1 and $\beta$ at 10. We averaged the results across three different seeds to ensure reliability and reproducibility. The standard deviation could be not reported if negligible.
@@ 5 Results
5.1 Fine-tuning performance
First, we evaluate the effectiveness of fine-tuning the entire model after pre-training procedure. The results presented in Table 3 indicate that, among all the evaluated pre-training methods, MLEM consistently achieves the most favorable results across all datasets, except for the highly sparse PhysioNet dataset. Furthermore, it is noteworthy that solely supervised learning consistently produces inferior results.
@@ 5.2 Embeddings probing
We conducted a comprehensive analysis of the quality of embeddings using both linear and non-linear probing techniques, results are presented in Table 4.
Our evaluation reveals that neither the contrastive nor the generative approach consistently outperforms the other. However, by effectively integrating the advantages of both contrastive and generative techniques, MLEM model demonstrates superior performance across most datasets on average. Notably, when MLEM is not the top performer, it is often ranked second, highlighting its versatility.
Table 5: Average anisotropy and intrinsic dimension across datasets
@@
Table 6: Performance change in linear probing performance after shuffling events order, higher is better
@@
5.3 Anisotropy and Intrinsic Dimension
We conducted an evaluation of anisotropy and the intrinsic dimensions of embeddings from various pre-training strategies, as detailed in 4.2.
Results presented in Table 6 show that MLEM has the highest intrinsic dimension and the lowest anisotropy, on average, across all datasets. This finding may indicate that our pre-training strategy effectively enables the embeddings to encapsulate a greater amount of information and ensures their uniform distribution in space.
However, we found no correlation between intrinsic dimension or anisotropy and downstream performance in our domain, as illustrated in correlation plot 4(b). While our findings align with the most recent study by Dubois et al. (2024) regarding anisotropy, they differ in that they observed a correlation between intrinsic dimension and performance.
Figure 3: The t-SNE visualizations showcase embeddings resulting from various pre-training strategies. The left portion displays the embeddings for the Age dataset, while the bottom row illustrates those for the TaoBao dataset. Each point represents a sequence $S_{i}$ from a given dataset, colored accordingly to the corresponding attribute $y_{i}$. For Age, there are 4 classes and 2 classes for TaoBao
5.4 Visualization
6 Embedding Robustness
Our study also explores the robustness of model embeddings to specific types of perturbations. We focus on examining the impact of dropout and the shuffling of events within one sequence on the linear probing performance.
We assess robustness with classification or regression tasks. In each scenario, we assess either the class of the whole sequence or the real value describing the whole sequence (pendulum length). We refer to this assessment as global and assume that this attribute does not change over the entire sequence (e.g., user age group). Attributes that change over time are referred to as local (e.g., amount of payment or next purchase category).
6.1 Sensitivity to permutation
We randomly permuted order of events along time axis and analyzed performance of pre-trained models. Models were trained without permutation. The results, presented in Table 6, show that the sequence order has only a small effect on most of the results. This suggests that current models may be employing a ’Bag-of-Words’ approach, treating the events as a set rather than considering the sequential nature of the data when determining certain global attributes. Additionally, we observed that the dataset’s degradation is greater when it has fewer categorical features.
Despite the small degradation, it is still present, highlighting the need for modeling temporal relations. Furthermore, all the models demonstrate the ability to encode sequential data, as evidenced by a decline in performance on the pendulum dataset where the time component is crucial.
The assessment of the effect of order permutation on the ability to model local attributes is similar to the TPP task, and we leave it for the future work.
6.2 Sensitivity to data omission
To investigate the effects of data omission, we applied a random dropout strategy, removing entire events from a sequence along the time axis, with varying probabilities $p_{dropout}\in\{0.1,0.3,0.5,0.7\}$. We report the average percentage decrease in linear probing performance, depicted in Figure 4(a).
We notice that MLEM performance slightly decreases at $0.1$ dropout probability and more so at higher levels. This indicates that MLEM is more sensitive to dropout and requires the presence of all the samples rather than depending on a couple specific samples. This property can be used in a sensitivity analysis or related applications.
Additionally, the generative modeling embeddings exhibited an intriguing behavior: they not only resisted the decline, but actually surpassed their original performance at a dropout rate of 0.5. We believe that such effect show high sensitivity to the presence of noisy or out-of-distribution events. Nevertheless, to draw reliable conclusions, further work is required.
(a) Mean percentage change averaged across datasets. The X-axis represents the dropout probability, the Y-axis indicates the mean change in the linear probing, compared to the metric with no dropout. Error bars reflect 1 sigma deviation.
(b) Correlation plots for anisotropy and intrinsic dimension with normalized performance.
@@ 7 Computational complexity
To obtain the MLEM encoder, we need to train two models, which is not desirable. However, to demonstrate the practicality of our method, we compare MLEM with SeqNas Udovichenko et al. (2024) in terms of performance and computational efficiency.
Table 7: Comparison of downstream metrics and computational costs for training MLEM versus SeqNas.
@@
SeqNas employs an architectural search to identify the most optimal supervised architecture for the task at hand. We posit that SeqNas represents an upper bound for downstream metrics. Table 9 presents results from the two largest datasets, examined by us, demonstrating that MLEM achieves performance near the upper bound set by SeqNas, but with significantly lower computational cost. Detailed compute budget reported in table 9
8 Related work
Generative methods for pre-training and representation learning have seen notable development in recent years in both NLP Radford and Narasimhan (2018); Devlin et al. (2019); Raffel et al. (2023) and CV He et al. (2021); Assran et al. (2023); Bachmann et al. (2022). In the EvS domain, the sole study by Lin et al. (2022) investigates various generative models for TPP tasks. Based on their findings, we have considered their best performing setting to serve as a generative model for our study, with slight modifications to fit our tasks.
Contrastive methods are a widely recognized in He et al. (2020); Grill et al. (2020). These methods typically draw closer the embeddings of variations of the same object and distance them from those of different objects. In CoLES Babaev et al. (2022) authors study contrastive learning for EvS. In their work, the positive samples consist of subsequences derived from a single user’s events, while the negative samples are obtained from events of different users. We utilize this method as best performing along contrastive approaches.
Hybrid methods. Our work falls into the category of hybrid self-supervised approaches Qi et al. (2023); Yang et al. (2023); Lazarow et al. (2017) Few studies have merged generative and contrastive methods, typically focusing on loss combination. DINOv2 Oquab et al. (2023) updates the traditional contrastive loss from DINO Caron et al. (2021) with iBOT’s Zhou et al. (2022) reconstruction loss through Masked Image Modeling. Similarly, GCRL Kim et al. (2021) combines these losses via a weighted sum, significantly outperforming models that are purely generative or contrastive. We follow the same methodology to form the hybrid baseline method with simply summing contrastive and generative losses.
Supervised learning is commonly used for the EvS classification task. The most comprehensive recent paper SeqNAS Udovichenko et al. (2024) adopts Neural Architecture Search to identify optimal architectures without strong prior assumptions about the underlying process. A key finding is the effectiveness of Recurrent Neural Networks (RNNs) in modeling EvS, leading us to incorporate RNN encoder in our model. Furthermore, the authors propose benchmark datasets for EvS classification. We selected the largest datasets from the benchmark, as self-supervised methods require substantial data volumes.
9 Limitations
Physionet underperformane. We observe consistent underperformance of methods that include generative approach on the PhysioNet dataset. This could be attributed to the high rate of missing values, which complicates the generation of accurate predictions for subsequent steps. Furthermore, the relatively small size of the dataset may be a contributing factor, given that generative modeling is generally dependent on large amounts of data.
Compute Resources. One limitation of our model is the requirement for a higher computational budget compared to previous methods. However, in the domain of event sequences, the primary challenge is typically achieving the highest metric rather than optimizing for computational efficiency.
Datasets and Models. We selected the most relevant datasets to cover a broad spectrum of tasks and chose the best-performing models from previous studies to represent SSL methods. Nonetheless, the comparison could be improved by incorporating a wider variety of datasets and model architectures.
Scalability. Although we aimed to ensure a fair comparison, the results may vary with different scales of data or models. However, observed trend show that more data leads to better results on MLEM and generative method.
@@ 10 Conclusion
To our knowledge, our study is the first to apply generative modeling to event sequences for representation learning and pre-training. We found that neither generative nor contrastive pre-training definitively outperforms the other.
In response, we developed the MLEM method, which creatively combines contrastive learning with generative modeling. Inspired by recent advances in multimodal research, MLEM treats these approaches as separate modalities and consistently outperforms existing models on most datasets across both primary and secondary objectives. This finding suggests a promising new direction for enhancing self-supervised learning across various domains.
Additionally, our research provides several key insights. Firstly, the best current self-supervised models do not utilize the sequential nature of data effectively. We also discovered that random event deletion does not adversely affect performance; in some cases, it actually enhances it. Secondly, unlike findings in other domains (Dubois et al., 2024; Ait-Saada and Nadif, 2023), we observed no correlation between anisotropy or intrinsic dimension and downstream performance.